# ECAMP: Entity-centered Context-aware Medical Vision Language Pre-training

### v3

@@ 3.3 Context-guided Super-resolution
Figure 3: The framework of multi-scale context fusion. The addition operation (highlighted in cyan) concurrently fuses the text feature, local aggregated vision feature, and global vision feature.
@@ 3.4 Multi-scale Context Fusion
Different from MRM [63] which directly adds the vision feature $f_{v}$ to the text embedding $E_{t}$, we first apply a self-attention (SA) for an initial comprehension of text embedding $f_{t}=\mathrm{SA}(E_{t})$. In order to simultaneously attain representative global and local features for multi-scale downstream tasks, we leverage patch-wise vision features $f_{v}$ as local vision features $f_{v}^{l}$ and use global average pooling (GAP) to compute global vision feature $f_{v}^{g}=\mathrm{GAP}(f_{v}^{l})$. Next, we aggregate local vision context $f_{v}^{l}$ with text feature $f_{t}$ by cross-attention (CA) to generate local aggregated feature $f_{a}^{l}=\mathrm{CA}(f_{v}^{l},f_{t})$, and project the $f_{v}^{g}$ by a linear projector to obtain global feature $f_{a}^{g}$. Finally, we integrate text feature $f_{t}$, local aggregated feature $f_{a}^{l}$, and duplicate global feature $f_{a}^{g}$ by addition.
@@
MedIM [56] - - - 88.9 89.3 89.7 - - - 77.2 90.3 93.6
ECAMP (Ours) 81.0 83.8 85.1 89.3 89.4 89.7 91.4 92.9 93.5 86.3 95.0 96.0
4 Experiment
4.1 Pre-training Dataset
MIMIC-CXR 2.0.0 [19, 20] contains 377,110 chest radiographs and 227,835 corresponding reports. Each report may be associated with one or more images captured from frontal, back, and lateral views. We extract detailed descriptions of medical diseases by focusing on the impression and finding sections in the free-text reports.
4.2 Downstream Tasks Datasets
ChestX-ray14 (CXR14) [51] comprises 112,120 frontal-view chest radiographs, and introduces a multi-label classification task involving 14 common chest pathologies. The dataset is officially split into training, validation, and testing subsets with a ratio of 70%/10%/20%.
CheXpert (CXP) [18] includes 223,648 chest radiographs from a frontal or lateral view, which categorizes each image into 5 diseases: atelectasis, cardiomegaly, consolidation, edema, and pleural effusion. Aligned with [61, 17], we designate the expert-labeled official validation set as the test data. We randomly chose 5,000 radiographs from the training data as the validation set. The split for training/validation/testing comprises 218,414/5,000/234 images.
RSNA Pneumonia [43] is composed of 29,700 frontal-view chest radiographs, each accompanied by pneumonia opacity bounding boxes if pneumonia is present in the image. Four tasks are conducted on this dataset: fine-tune classification, linear classification, semantic segmentation, and object detection. For fine-tune and linear classification tasks, we adhere to the official data partitioning, allocating 25,184, 1,500, and 3,000 images for the training, validation, and test sets, respectively. In the context of semantic segmentation and object detection, our approach aligns with the methodology proposed in [49], wherein the original training set is randomly divided into subsets of 16,010, 5,337, and 5,337 samples for training, validation, and testing purposes, respectively.
COVIDx CXR-3 [50] comprises 29,986 chest radiographs. This dataset presents a multi-class classification task, categorizing each image into COVID-19, non-COVID pneumonia, or normal. In line with [49, 48, 5], version 6 of this dataset is used. The official validation set is considered as our test data, and simultaneously, 10% of the original training set is randomly selected as validation data.
SIIM-ACR Pneumothorax [44] introduces the task of pneumothorax segmentation for evaluating the performance of segmentation models. This dataset includes 12,047 frontal-view chest radiographs with meticulously manually labeled pneumothorax masks. In accordance with [17], we partition the dataset into training/validation/test sets with a ratio of 70%/15%/15%.
@@ 4.2 Downstream Tasks Datasets
MedIM [56] - 63.5 81.3 - - - - - -
ECAMP (Ours) 67.3 70.4 84.5 77.1 78.6 79.5 19.8 25.8 31.0
@@ 4.4 Implementation Details
We utilize an 8-layer ViT-B/16 [12] as the backbone for the image encoder and a 4-layer ViT-B/16 for the image decoder. Additionally, we employ BERT [11] as the text encoder. If not otherwise stated, we resize images to 224$\times$224 as input, except for super-resolution supervision we retain the 448$\times$448 resolution images. Texts are padded or truncated to a fixed length of 256 tokens. The latent dimension, denoted as D, is set to 768. The implementation is based on PyTorch [38]. For pre-training, our framework runs for 115 epochs on 4 NVIDIA A100 GPUs, taking approximately 20 hours. The training batch size is set to 256 for each GPU, and we utilize a gradient accumulation step of 8. Consequently, the effective batch size reaches 8,192. We opt for AdamW [33] as our default optimizer, with a learning rate of 1.5e-4 and weight decay of 0.05. The values for $\beta_{1}$ and $\beta_{2}$ are set to 0.9 and 0.95, respectively. The mean squared error (MSE) loss is employed for masked image modeling and super-resolution, while cross-entropy loss is used for masked language modeling.
@@
For the LLM in Eq. 1, we use GPT-3.5-turbo API [36] with setting temperature to 0 to distill 10,000 reports and fine-tune a Vicuna model [8] to distill the rest reports.
@@ 4.5 Baselines
We conduct a comprehensive comparative analysis of ECAMP with numerous SOTA Med-VLP approaches pre-trained solely on MIMIC-CXR using ResNet-50 [16] (‘CNN-Based’) or ViT-B/16 [12] (‘ViT-Based’) as backbone, including MAE [15] trained on ImageNet [21] (denoted as ‘ImageNet init’), GLoRIA [17], REFERS [62], SAT [31], PRIOR [7], MedKLIP [54], MGCA [49], KAD [60], MedIM [56], and MRM [63]. Notably, Med-UniC [48] uses external multilingual datasets. In addition, we pre-train MAE [15] (denoted as ‘MAE’) and GLoRIA [17] with ViT-B/16 on MIMIC-CXR [19] using official source code. For MGCA, MRM, SAT, PRIOR, KAD, and MedKLIP whose official model are released, we fine-tune their models on the public downstream datasets in case there is no corresponding performance metrics in referenced papers. For MedIM and Med-UniC, whose official models are absent, we directly copy the performances from their papers for comparison.
Table 4: Ablation study of five simple-yet-effective components (context distillation (CD), descriptors masking (DM), re-balanced factor (RF), context-guided super-resolution (SR), and MSCF) proposed in ECAMP on multi-scale downstream tasks including fine-tune classification (CXR14 [51] ), linear classification (COVIDx [50]), semantic segmentation (SIIM [44]), and object detection (RSNA [43]). All of the downstream tasks are trained with 1% training data.
CD DM RF SR MSCF CXR14 (AUC) COVIDx (ACC) SIIM (Dice) RSNA (mAP)
Global Local
$\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ 79.5 78.0 65.4 4.1
$\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ 80.2 81.5 65.9 12.2
$\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ 81.1 82.3 66.5 16.1
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ 81.5 83.0 66.6 14.9
$\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 81.3 82.5 66.9 15.6
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 81.4 83.0 66.4 16.8
$\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 81.1 84.8 67.1 18.9
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 81.8 85.0 65.7 12.6
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 81.4 83.3 67.5 20.4
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 81.9 86.3 67.3 19.8
@@ 5.1 Results on Classification
Table 5: Ablation study on parameter $\lambda_{neg}$ proposed in Eq. 3 on fine-tuning (CXR14 [51]) and linear (COVIDx [50]) classification with different ratios of training data (1%, 10%).
$\lambda_{neg}$ CXR14 (AUC) COVIDx (ACC)
1% 10% 1% 10%
0.025 81.7 84.5 85.0 93.3
0.05 81.9 85.1 86.3 95.0
0.1 81.8 84.6 80.5 94.8
0.25 81.5 84.2 79.5 94.0
0.5 81.4 84.2 82.5 91.5
It is worth noting that, on the ChestX-ray14 [51], which encompasses the most of testing data and disease labels, ECAMP surpasses the strongest counterpart by 2.5% using only 1% labeled data. For a comparison of 14 diseases, refer to Appendix.D. Surprisingly, on the COVIDx with a novel disease “COVID-19”, which is not present in MIMIC-CXR [19], a great performance gap can be found in Table 1 between ECAMP and SOTA methods. This validates the superiority of the representation learned by our ECAMP. Interestingly, we find the cross-modality reconstruction-based methods MRM [63] and our ECAMP outperform the contrastive-based methods like MGCA [49] and GloRIA [17]. Moreover, Table 2 presents the results of linear-probe classification. ECAMP exhibits consistently improved performance compared with SOTA methods on CXR14, RSNA and COVIDx datasets.

### v4

@@ 3.4 Multi-scale Context Fusion
Different from MRM [73] which directly adds the vision feature $f_{v}$ to the text embedding $E_{t}$, we first apply a self-attention (SA) for an initial comprehension of text embedding $f_{t}=\mathrm{SA}(E_{t})$. In order to simultaneously attain representative global and local features for multi-scale downstream tasks, we leverage patch-wise vision features $f_{v}$ as local vision features $f_{v}^{l}$ and use global average pooling (GAP) to compute global vision feature $f_{v}^{g}=\mathrm{GAP}(f_{v}^{l})$. Next, we aggregate local vision context $f_{v}^{l}$ with text feature $f_{t}$ by cross-attention (CA) to generate local aggregated feature $f_{a}^{l}=\mathrm{CA}(f_{v}^{l},f_{t})$, and project the $f_{v}^{g}$ by a linear projector to obtain global feature $f_{a}^{g}$. Finally, we integrate text feature $f_{t}$, local aggregated feature $f_{a}^{l}$, and duplicate global feature $f_{a}^{g}$ by addition. During the pre-training stage, the only supervision which contains diagnosis information in our framework comes from MLM loss (Eq. 4). So, as shown in Fig. 4, the gradient is propagated to local vision feature $f_{v}^{l}$ and global vision feature $f_{v}^{g}$ via addition and duplication simultaneously, and train the multi-scale features.
@@ 4 Experiment
MedIM-C [64] - - - 88.9 89.3 89.7 - - - 77.2 90.3 93.6
MLIP# [34] - - - 89.0 89.4 90.0 89.3 90.0 90.8 75.3 86.3 92.5
ECAMP (Ours) 81.0 83.8 85.1 89.3 89.5 89.7 91.4 92.9 93.5 86.3 95.0 96.0
@@
MedIM-C [64] - 63.5 81.3 - - - - - -
MedIM [65] - 64.2 82.0 - - - - - -
ECAMP (Ours) 67.3 70.8 84.5 77.1 78.6 79.5 19.8 28.0 31.6
4.1 Pre-training Dataset
MIMIC-CXR 2.0.0  [^4] https://physionet.org/content/mimic-cxr/2.0.0/ [21, 22] contains 377,110 chest radiographs and 227,835 corresponding reports. Each report may be associated with one or more images captured from frontal, back, and lateral views. We extract detailed descriptions of medical diseases by focusing on the impression and finding sections in the free-text reports.
FFA-IR  [^5] https://physionet.org/content/ffa-ir-medical-report/1.0.0/ [29] includes 10,790 reports along with 1,048,584 Fundus Fluorescein Angiography (FFA) images from clinical practice. It comprises explainable annotations, based on a schema of 46 categories of lesions; and it is bilingual, providing both English and Chinese reports for each case. We select English reports and paired FFA images as pre-training data. This setup is used to evaluate the performance of pre-training frameworks on fundoscopy images.
4.2 Downstream Tasks Datasets
ChestX-ray14  [^6] https://nihcc.app.box.com/v/ChestXray-NIHCC [58] comprises 112,120 frontal-view chest radiographs, and introduces a multi-label classification task involving 14 common chest pathologies. The dataset is officially split into training, validation, and testing subsets with a ratio of 70%/10%/20%.
CheXpert  [^7] https://stanfordmlgroup.github.io/competitions/chexpert/ [18] includes 223,648 chest radiographs from a frontal or lateral view, which categorizes each image into 5 diseases: atelectasis, cardiomegaly, consolidation, edema, and pleural effusion. Aligned with [71, 17], we designate the expert-labeled official validation set as the test data. We randomly choose 5,000 radiographs from the training data as the validation set. The split for training/validation/testing comprises 218,414/5,000/234 images.
RSNA Pneumonia (RSNA)  [^8] https://www.rsna.org/rsnai/ai-image-challenge/rsna-pneumonia-detection-challenge-2018 [49] is composed of 29,700 frontal-view chest radiographs, each accompanied by pneumonia opacity bounding boxes if pneumonia is present in the image. Four tasks are conducted on this dataset: fine-tune classification, linear classification, semantic segmentation, and object detection. For fine-tune and linear classification tasks, we adhere to the official data partitioning, allocating 25,184, 1,500, and 3,000 images for the training, validation, and test sets, respectively. In the context of semantic segmentation and object detection, our approach aligns with the methodology proposed in [56], wherein the original training set is randomly divided into subsets of 16,010/5,337/5,337 samples for training/validation/testing purposes, respectively.
COVIDx CXR-3 (COVIDx)  [^9] https://www.kaggle.com/datasets/andyczhao/covidx-cxr2/versions/7 [57] comprises 29,986 chest radiographs. This dataset presents a multi-class classification task, categorizing each image into COVID-19, non-COVID pneumonia, or normal. In line with [56, 55, 5], version 6 of this dataset is used. The official validation set is considered as our test data, and simultaneously, 10% of the original training set is randomly selected as validation data.
SIIM-ACR Pneumothorax (SIIM)  [^10] https://www.kaggle.com/c/siim-acr-pneumothorax-segmentation [51] introduces the task of pneumothorax segmentation for evaluating the performance of segmentation models. This dataset includes 12,047 frontal-view chest radiographs with meticulously manually labeled pneumothorax masks. In accordance with [17], we partition the dataset into training/validation/test sets with a ratio of 70%/15%/15%.
ODIR-5K  [^11] https://odir2019.grandchallenge.org/ consists of 5,000 patient records, each containing color fundus photographs of both the left and right eyes, along with corresponding diagnostic keywords provided by clinicians. These keywords are categorized into eight distinct classes. We utilize the official test set as the test data. Additionally, 639 images are randomly selected as validation data, resulting in a data split of 70% for training, 10% for validation, and 20% for testing.
APTOS-2019  [^12] https://www.kaggle.com/datasets/mariaherrerot/aptos2019 [23] presents a multi-class classification task for determining the severity level of diabetic retinopathy (DR) in fundus images. The dataset comprises 3,662 labeled images, with DR severity levels ranging from 0 to 4. The dataset is randomly partitioned into training, validation, and test sets with a split ratio of 70%/10%/20%, respectively.
MuReD  [^13] https://data.mendeley.com/datasets/pc4mb3h8hz/1 [47] consists of 2,451 multi-label retinal fundus images covering 20 categories. We randomly select 222 images from the official training set to serve as our validation data. The official test set is used for testing data. The data is split into 70% for training, 10% for validation, and 20% for testing.
RIGA  [^14] https://deepblue.lib.umich.edu/data/concern/data_sets/3b591905z [20] introduces the task of retinal cup and disc segmentation, containing a total of 750 color fundus images from three sources. We adopt the official test set, with the dataset split into 590 images for training, 95 for validation, and 95 for testing.
@@ 4.4 Implementation Details
We utilize an 8-layer ViT-B/16 [12] as the backbone for the image encoder and a 4-layer ViT-B/16 for the image decoder. Additionally, we employ BERT [11] as the text encoder. If not otherwise stated, we resize images to 224$\times$224 as input, except for super-resolution supervision we retain the 448$\times$448 resolution images. Texts are padded or truncated to a fixed length of 256 tokens. The latent dimension, denoted as D, is set to 768. The implementation is based on PyTorch [43]. For pre-training on MIMIC-CXR [21, 22], our framework runs for 120 epochs on 4 A100 GPUs, taking approximately 20 hours. The training batch size is set to 256 for each GPU, and we utilize a gradient accumulation step of 8. Consequently, the effective batch size reaches 8,192. For pre-training on FFA-IR [29], our framework runs for 60 epochs on 4 A100 GPUs, taking approximately 15 hours. Other settings are the same with the pre-training stage on CXRs. We opt for AdamW [38] as our default optimizer, with a learning rate of 1.5e-4 and weight decay of 0.05. The values for $\beta_{1}$ and $\beta_{2}$ are set to 0.9 and 0.95. The mean squared error (MSE) loss is employed for masked image modeling and super-resolution, while cross-entropy loss is used for masked language modeling.
@@
For the LLM in Eq. 1, we use GPT-3.5-turbo API [41] with temperature setting to 0 to distill 10,000 reports and fine-tune a Vicuna model [8] to distill the rest of the reports.
Table 7: Ablation study of five simple-yet-effective components (context distillation (CD), descriptors masking (DM), re-balancing factor (RF), context-guided super-resolution (SR), andMSCF) proposed in ECAMP on multi-scale downstream tasks including fine-tune classification (ChestX-ray14 [58] ), linear classification (COVIDx [57]), semantic segmentation (SIIM [51]), and object detection (RSNA [49]). All of the downstream tasks are trained with 1%, 10% and 100% of training data.
CD DM RF SR MSCF ChestX-ray14 (AUC) COVIDx (ACC) SIIM (Dice) RSNA (mAP)
Global Local 1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100%
$\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ 79.5 83.9 86.0 78.0 87.3 92.3 65.4 68.6 80.7 4.1 20.6 26.0
$\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ 80.2 84.2 86.1 81.5 89.0 89.5 65.9 69.2 81.2 12.2 22.4 26.6
$\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ 80.8 84.4 86.2 82.1 90.3 92.8 66.0 69.4 82.4 13.9 23.3 27.3
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ 81.5 84.7 86.5 83.0 92.3 95.0 66.6 69.9 83.6 16.1 24.5 29.1
$\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 81.3 84.6 86.3 82.5 92.5 95.0 66.9 69.9 84.1 15.6 25.2 29.7
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 81.4 84.5 86.2 83.0 94.0 95.3 66.4 70.0 83.4 16.8 25.7 30.5
$\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 80.8 84.1 86.2 82.8 91.8 92.5 66.6 69.8 84.0 16.9 26.5 30.3
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 81.8 85.0 86.4 85.0 92.5 94.5 65.7 69.5 83.6 12.6 21.7 28.6
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 81.4 84.7 86.1 83.3 88.0 90.8 67.5 70.7 84.3 20.4 28.2 31.7
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 81.9 85.1 86.7 86.3 95.0 96.0 67.3 70.8 84.5 19.8 28.0 31.6
Table 8: Ablation study on the impact of two critical components: Entity Selection Strategy and Text Input Composition. We compare three entity selection strategies: (1) Top-20 highest-frequency clinical entities, (2) Top-10% frequency-percentile cutoff, and (3) our proposed method leveraging the Top-44 most frequent entities. For text input composition, we assess four paradigms: (1) original radiology reports alone (Origin-alone), (2) ChatGPT-distilled reports alone (Dis-alone), (3) hybrid inputs combining original reports with named entity recognition (NER)-based distilled reports (Hybrid-NER), and (4) hybrid inputs combining original reports with ChatGPT-distilled reports (Hybrid-GPT). These configurations are evaluated on multi-scale downstream tasks including fine-tune classification (ChestX-ray14 [58]), linear classification (COVIDx [57]), semantic segmentation (SIIM [51]), and object detection (RSNA [49]). All of the downstream tasks are trained with 1%, 10% and 100% of training data.
Entity Text Input ChestX-ray14 (AUC) COVIDx (ACC) SIIM (Dice) RSNA (mAP)
Selection Composition 1% 10% 100% 1% 10% 100% 1% 10% 100% 1% 10% 100%
Top-20 Hybrid-GPT 81.2 84.7 86.4 84.8 92.8 94.5 66.7 70.6 82.5 13.3 25.3 26.9
Top-10% Hybrid-GPT 80.5 84.0 86.3 83.5 90.5 93.3 66.6 70.6 83.1 17.1 25.6 28.0
Top-44 Origin-alone 81.3 84.6 86.3 82.5 92.5 95.0 66.9 69.9 84.1 15.6 25.2 29.7
Top-44 Dis-alone 80.8 84.2 86.2 82.0 91.0 94.5 67.1 69.6 84.3 15.3 25.4 29.4
Top-44 Hybrid-NER 81.7 84.9 86.6 84.0 93.8 95.3 66.9 70.1 84.2 16.2 25.9 30.4
Top-44 Hybrid-GPT 81.9 85.1 86.7 86.3 95.0 96.0 67.3 70.8 84.5 19.8 28.0 31.6
@@ 4.5 Baselines
We conduct a comprehensive comparative analysis of ECAMP with numerous SOTA Med-VLP approaches pre-trained solely on MIMIC-CXR using ResNet-50 [16] (‘CNN-Based’) or ViT-B/16 [12] (‘ViT-Based’) as backbone, including MAE [15] trained on ImageNet [24] (denoted as ‘ImageNet init’), GLoRIA [17], REFERS [72], SAT [36], PRIOR [7], MedKLIP [61], MGCA [56], KAD [70], MedIM-C [64], MRM [73], MedIM [65] and MLIP [34]. Notably, Med-UniC [55] uses external multilingual datasets. In addition, for fair comparison, we pre-train MAE [15] (denoted as ‘MAE’) and GLoRIA [17] with ViT-B/16 on MIMIC-CXR [21] using official source code. For MGCA, MRM, SAT, PRIOR, KAD, and MedKLIP whose official model are released, we fine-tune their models on the public downstream datasets in case there is no corresponding performance metrics in referenced papers. For MedIM-C, Med-UniC, MedIM and MLIP, whose official models are absent, we directly copy the performances from their papers for comparison. For extensive experiments on FFA images, we reproduce MAE [15], GLoRIA [17] and MRM [73] using their official code pre-trained on FFA-IR [29] dataset, which is under the same settings as our ECAMP.
Fig. 5: T-SNE visualizations of ECAMP and MRM [73] in the fine-tuning classification task on COVIDx [57] dataset. The left panel represents ECAMP, while the right panel shows MRM. The blue points correspond to images labeled as “no pneumonia”, the green points represent image features related to “pneumonia”, and the red points indicate “COVID-19” features.
@@ 5.1 Results on Classification
Table 9: Ablation study on parameter $\lambda_{neg}$ proposed in Eq. 3 on fine-tuning (ChestX-ray14 [58]) and linear (COVIDx [57]) classification with different ratios of training data (1%, 10%, and 100%).
$\lambda_{neg}$ ChestX-ray14 (AUC) COVIDx (ACC)
1% 10% 100% 1% 10% 100%
0.025 81.7 84.5 86.5 85.0 93.3 94.8
0.05 81.9 85.1 86.7 86.3 95.0 96.0
0.1 81.8 84.6 86.4 80.5 94.8 95.3
0.25 81.5 84.2 86.2 79.5 94.0 94.5
0.5 81.4 84.2 86.2 82.5 91.5 91.8
It is worth noting that, on the ChestX-ray14 [58], which encompasses the most of testing data and disease labels, ECAMP surpasses the strongest counterpart by 2.5% using only 1% labeled data. Surprisingly, on the COVIDx with a novel disease “COVID-19”, which is not present in MIMIC-CXR [21], a great performance gap can be found in Table 4 between ECAMP and SOTA methods. COVIDx dataset includes the RSNA pneumonia dataset as well as a collection of chest X-rays featuring COVID-19 cases, which is similar to pneumonia. So, the pre-trained method has to learn different textures of pneumonia to further classify the COVID-19 of normal pneumonia. As shown in Fig. 5, we use t-SNE to visualize the features of our ECAMP and MRM. It is evident that our model demonstrates a better separation of features across “no pneumonia”; “pneumonia”, and “COVID-19”, which supports the superiority of pre-trained features of our ECAMP method and possibly enables better generalizability to novel class. This validates the superiority of the representation learned by our ECAMP. Interestingly, we find the cross-modality reconstruction-based methods MRM [73] and our ECAMP outperform the contrastive-based methods like MGCA [56] and GloRIA [17]. Moreover, Table 5 presents the results of linear-probe classification. ECAMP exhibits consistently improved performance compared with SOTA methods on ChestX-ray14, RSNA and COVIDx datasets. To evaluate the generalizability and adaptability of ECAMP, we extend the pre-training framework to Fundus Fluorescein Angiography (FFA) images and corresponding textual reports. Comprehensive experiments on fine-tuning classification is presented in Table 10. We perform fine-tuning for both multi-label and multi-class classification. As the table shows, our proposed ECAMP method substantially outperforms all state-of-the-art ViT-based medical vision-language pre-training approaches, even when applied beyond the CXR image domain. Notably, on the ODIR-5K dataset, which includes labels representing a broad spectrum of fundus diseases, ECAMP achieves a 5% improvement over the strongest competing method, despite utilizing only 1% of the labeled data.
[…]
Method ODIR-5K (AUC) APTOS (ACC) MuRed (AUC)
1% 10% 100% 1% 10% 100% 10% 100%
Random init 54.1 57.5 65.7 59.7 63.4 68.5 51.1 66.5
ImageNet init 62.6 64.0 77.4 68.3 71.7 75.0 52.9 70.7
ViT-based
MAE[15] 63.8 72.9 82.4 71.8 75.0 79.9 68.8 82.9
GLoRIA[17] 67.4 74.1 87.2 72.3 74.9 81.0 81.0 93.5
MRM[73] 66.8 78.3 88.1 73.9 74.3 80.9 82.1 93.0
ECAMP (Ours) 71.8 81.2 90.0 75.9 76.9 83.2 85.6 94.5